OpenStudy (madgirlwithabluebox):

will fan and medal. Suppose a parabola has vertex (–4, 7) and also passes through the point (–3, 8) How would you write this in vertex form?

OpenStudy (anonymous):

do u know the standard equation of vetex form??/

OpenStudy (anonymous):

it is : y=a(x-h)^2 + k where (h,k) is vertex...

OpenStudy (anonymous):

now u cn solve?

OpenStudy (madgirlwithabluebox):

y=(x+4)^2 +7 ?

OpenStudy (anonymous):

correct..:)

OpenStudy (madgirlwithabluebox):

Can you help me with 2 more?

OpenStudy (anonymous):

yeah sure

OpenStudy (madgirlwithabluebox):

Write the expressions in factored form. 9x^2-4 4x^2+11x+6 6x^2+4x x^2+13x+42

OpenStudy (anonymous):

we have three basic forms...: \[a ^{2} - b ^{2} = (a+b)(a-b)\] \[(a+b)^{2} = a ^{2} +2 a b + b ^{2}\] \[(a-b)^{2}=a ^{2}-2ab+b ^{2}\]

OpenStudy (anonymous):

so our first question is of first form.. solve and tell..

OpenStudy (anonymous):

for second question its splitting of middle term... do u know how to do it?

OpenStudy (anonymous):

in third question... see we can take something out common... so take common out.. put else in bracket....

OpenStudy (anonymous):

and fourth is again splitting...

OpenStudy (madgirlwithabluebox):

i have no idea where to start though

OpenStudy (anonymous):

okk... for first question: we can write as (3x)^2 - (2)^2

OpenStudy (anonymous):

understood?

OpenStudy (madgirlwithabluebox):

not really?

OpenStudy (madgirlwithabluebox):

so 6x^2+4x would be 2x(3x + 2)

OpenStudy (anonymous):

right ,....